OpenStudy (anonymous):

An isosceles triangle has an area of 24 cm^2, and the angle between the two equal sides is 5π/6. What is the length of the two equal sides? (Round your answer to one decimal place.)

OpenStudy (anonymous):

9.79

OpenStudy (anonymous):

is that right ans.

OpenStudy (campbell_st):

use the trig ratio for area of a triangle... and call the equal sides a  then you say \[24 = \frac{1}{2} \times a \times a \times \sin(\frac{5\pi}{6})\] so simplifying things a little \[48 = a^2 \times \sin(\frac{5\pi}{6})\] now just solve for a, and only use the positive value.